Whats the annual interest rate, to the nearest 10th of a percent, for $1839 to grow to $25187 if the interest is compunded yearly for 38 years
Answer by Nate(3500)

P(a) = a(1 + r)^t
25187 = 1839(1 + r)^38
(25187/1839)^(1/38) = (1 + r)
(25187/1839)^(1/38) - 1 = r
0.07129823 = r
About 7%